### v1

# A candidate quadruple AGN system at

11email: eherwig@mpa-garching.mpg.de 22institutetext: Max Planck Institut für Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 33institutetext: International Gemini Observatory, NSF’s NOIRLab, 670 N A’ohoku Place, Hilo, Hawai’i 96720, USA
A candidate quadruple AGN system at
Multiple galaxies hosting active galactic nuclei (AGNs) at kpc separation from each other are exceedingly rare, and in fact, only one quadruple AGN is known so far. These extreme density peaks are expected to pinpoint protocluster environments and therefore be surrounded by large galaxy overdensities. In this letter, we present another quadruple AGN candidate at including two SDSS quasars at a separation of roughly 480 kpc. The brighter quasar is accompanied by two AGN candidates (a type 1 AGN and a likely type 2 quasar) at close ( kpc) separation identified through emission line ratios, line widths and high ionization lines like N v . The extended Ly emission associated with the close triple system is more modest in extent and brightness compared to similar multiple AGN systems and could be caused by ram-pressure stripping of the type-2 quasar host during infall into the central dark matter halo. The projected evolution of the system into a galaxy cluster with the AGN host galaxies forming the brightest cluster galaxy needs to be further tested by galaxy overdensity studies on large scales around the quadruple AGN candidate. If confirmed as a quadruple AGN with X-ray observations or rest-frame optical line ratios, this system would represent the second AGN quartet, the highest-redshift multiplet and the closest high-redshift triplet known.
Key Words.:
Galaxies: clusters: general – Galaxies: evolution – quasars: general1 Introduction
In recent years, efforts to identify the progenitors of galaxy clusters, so called protoclusters (e.g., Overzier 2016), have accelerated. Protoclusters in simulations are simply defined as structures eventually coalescing into a virialized galaxy group of mass at . Observationally, the identification of protoclusters is more complicated. High- quasars have long been believed to be signposts of peaks in the density distribution, but remain controversial as a method to identify protoclusters (Uchiyama et al., 2018; Habouzit et al., 2019). While they often do reside in overdensities (Fossati et al., 2021; García-Vergara et al., 2022), their typical halo masses of at are much lower than cluster masses (White et al., 2012; Trainor & Steidel, 2012; Husband et al., 2013), requiring substantial growth to form a galaxy cluster until .
While single bright quasars may pinpoint protocluster cores, physically associated multiple AGN systems could be even more reliable tracers of local overdensities (Onoue et al., 2018; Sandrinelli et al., 2018), albeit also not a firm indicator (Fukugita et al., 2004; Green et al., 2011), as a number of highly active, massive galaxies require ample supply of cold gas feeding supermassive black holes (SMBHs) and star formation. In the framework of hierarchical structure formation, such objects are most likely to reside in peaks of the density field populating the rare massive end of the halo mass function (Bhowmick et al., 2020). In the absence of dense environments, multiple AGN systems would be exceedingly rare due to their transient nature (Hennawi et al., 2006).
Interactions of three SMBHs are also predicted to be almost always necessary to form ultra-massive black holes with black hole masses (Hoffman et al., 2023), further highlighting their importance in cluster formation as the central brightest cluster galaxies commonly host ultra-massive black holes (e.g., Dullo et al. 2017; Mehrgan et al. 2019).
So far, only a few multiple AGN systems on kiloparsec scales are known. These objects are likely in an early stage of galaxy merger and might share a dark matter halo, but not their host galaxies. In contrast to that, sub-arcsecond triple AGNs trace local late-stage mergers with multiple cores in the same host galaxy (e.g., Liu et al. 2011; Kalfountzou et al. 2017; Pfeifle et al. 2019). This work concentrates on the former AGN tuples on kiloparsec scales, of which only five are known.
Specifically, Hennawi et al. (2015) report the discovery of a quasar quartet at the peak of cosmic noon, . The AGN are embedded in a massive Ly nebula extending by more than 300 kpc, indicative of a substantial reservoir of cool ( K) gas (). The AGN are all found within a sphere of projected radius ¡150 kpc. The first triple quasar to be identified also resides at a redshift of 2 within a projected distance of less than 50 kpc, consistent with the typical onset of AGN activity in galaxy mergers (Djorgovski et al., 2007). Another triplet has been detected at , consisting of a close pair with a companion at 200 kpc projected distance (Farina et al., 2013). The third triplet known is likewise embedded in a massive Ly nebula at and comprises two quasars and one faint AGN at projected distances below 100 kpc (Arrigoni Battaia et al., 2018). A fourth triplet was found in a sample of almost 20,000 AGN candidates at a redshift of 1.13 (Assef et al., 2018).
Here, we report on the discovery of a quadruple AGN candidate at consisting of a close triplet, contained within 20 kpc in projection, and a forth quasar at a separation of 478 kpc.
Throughout this work, we assume a flat CDM cosmology with , and (Planck Collaboration et al., 2020). Reported quasar magnitudes are in the AB system.
2 Observation and Data Analysis
One quasar in the system, SDSS J101254.73+033548.7 (hereinafter QSO1), has been targeted with the Multi Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) on the Very Large Telescope as part of a survey to study the environment of quasar pairs (Herwig et al., 2024). QSO1 is a known member of a quasar pair from the SDSS DR12 (Pâris et al., 2017). The projected distance to the second quasar in the pair SDSS J101251.06+033616.5 (hereinafter QSO2) is 478 kpc. For QSO2, only the SDSS spectrum is available. Here we briefly describe the observations and data reduction. More detailed descriptions of sample selection, observation and subsequent data analysis can be found in Herwig et al. (2024). The MUSE data were acquired in Wide Field Mode on March 14, 2018 (seeing of ), yielding a field of view (FoV) of (pixel scale of ) and targeting only QSO1 as the separation of the pair is larger than the MUSE FoV. Three exposures were taken with on-source time each, rotated by 90 degrees with respect to each other and covering a wavelength range of with a sampling of . The data have been reduced following Farina et al. (2019) and González Lobos et al. (2023). Specifically, we use the MUSE pipeline v 2.8.7 (Weilbacher et al., 2014, 2016, 2020) to perform bias- and sky-subtraction, flat fields, twilight and illumination correction and wavelength and flux calibration. To remove residual sky emission lines, we use the Zurich Atmospheric Purge (Soto et al., 2016). After masking gaps between the CCDs and correcting offsets between the exposures with custom routines, the three exposures are median combined to obtain the final data cube. To take into account correlated noise introduced during data processing, the variance cube produced by the pipeline is rescaled layer-by-layer to the RMS of a background region in the science cube, resulting in a mean pixel-wise noise level at the position of QSO1 of .
3 Results
3.1 Discovery of a candidate quadruple AGN
By collapsing the final MUSE cube along the wavelength axis, we obtain a continuum image of the field surrounding QSO1. In this image (r-band in Fig. 1, left), two additional continuum sources are visible in close proximity to QSO1. While similar geometries can be seen as a result of gravitational lensing of a single source, we can confidentily reject this possibility in this case (Appendix A). We instead identify both of these sources (hereinafter called AGNc1 and AGNc2) as AGN candidates as detailed in the following. To constrain their nature, we extract spectra from the MUSE cube within an aperture of 0.8 to minimize contamination from each other’s PSFs. To determine their positions, we first model the PSF of a star within the MUSE FoV with a 2D Moffat function and fit the model to the three continuum sources observed with MUSE (QSO1, AGNc1, AGNc2) by fixing the FWHM and leaving amplitude and position free to vary (Fig.1, middle panel). The peak of the Moffat function is used as center of the aperture for spectral extraction and is indicated with pink circles (crosses) in the middle (right) panel of Fig. 1. To reduce noise while retaining the signal, the spectra are smoothed by a 1D Gaussian kernel with pixels, corresponding to . Before smoothing the spectrum of AGNc2, a sky emission line at 6365.3 is masked as it would otherwise be convolved with the C iv line. The spectra are shown in Fig. 2, together with the SDSS spectrum of QSO2. The two QSOs and two AGN candidates are all detected in Ly, confirming that they reside at roughly the same redshift. QSO1 and AGNc1 both show a broad C iii] emission line, while in the spectrum of AGNc2, there is a tentative detection of a narrow C iii] line as well as a prominent narrow, blue-shifted C iv line. A He ii emission line is tentatively detected in all four objects. Within the aperture of used for spectral extraction, the -band magnitude for the three AGNs observed with MUSE is mag, mag and mag. In SDSS DR17 (Abdurro’uf et al., 2022), the -band magnitude of QSO2 is reported as . However, the SDSS fiber diameter of 3 is much larger than the aperture used in this work and consequently, the magnitude of QSO1 is reported to be much brighter in SDSS with . Between QSO1 and AGNc1, we measure an angular separation of 2.1, corresponding to the distance kpc at this redshift. QSO1 and AGNc2 are separated by 2.6 or 20 kpc, and QSO2 is 61.6 or 478 kpc away from QSO1. We obtain the systemic redshift by fitting the C iii] line with a double Gaussian model and correcting by the characteristic velocity shift of between the peak of the line complex and the systemic redshift for quasars (Shen et al., 2016). The value obtained for AGNc1 is likely overestimated as a redshift of implies a highly unusual Ly line shape that is completely dominated by the blue wing (Fig. 4). If this was the case, the He ii line might be misidentified and possibly not detected. Positions of the AGN candidates and projected distance between each other as well as redshift and velocity shifts are summarized in Table 1.
We fit the spectra as described in Appendix B and summarize the resulting line fluxes, widths and upper flux limits in Table 2. Black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived for C iv in Vestergaard & Peterson (2006) are and . The line width of AGNc2 is narrow with and consistent with a Type II AGN, while the spectrum of AGNc1 is dominated by broad lines of indicative of a Type I AGN.
At the position of QSO1, radio emission is detected with a flux density of in VLASS (Lacy et al., 2020) and in VLA FIRST (Becker et al., 1994), leading to a radio slope of (). Additionally using the UV continuum slope from the fit to the spectrum of QSO1 (Appendix B) of (), we can calculate the ratio (Kellermann et al., 1989) by extrapolating the rest-frame flux densities and obtain , classifying QSO1 as radio-loud. Although no additional radio emission is detected, we cannot exclude that AGNc1 and AGNc2 are radio-loud due to the shallow surveys depth and their dim UV continuum.
[…]
QSO1 | 10:12:54.74 | +03:35:48.93 | 3.166 0.001 | 0 | 0/ 0 | |
AGNc1 | 10:12:54.88 | +03:35:49.22 | 3.183 0.008 | -1590 766 | 2.1/16 | |
AGNc2 | 10:12:54.77 | +03:35:46.36 | 3.158 0.001 | 761 166 | 2.6/20 | |
[…]
3.1.1 Emission line diagnostics
A commonly adopted diagnostic diagram to distinguish AGNs from star forming galaxies (SFGs) in rest-frame UV uses the line ratios C iv /C iii] (C43), mainly influenced by the ionization parameter, and C iii]/He ii (C3He2), which is sensitive to metallicity and allows distinction between shocks and AGN photoionization (Feltre et al., 2016). Hirschmann et al. (2019) propose demarcation lines to distinguish between active and inactive galaxies, but current models are not capable of fully reproducing observed line ratios, leading to large impurities in the selection of AGNs and specifically faint AGNs. Classification is further complicated because certain types of galaxies would not appear in such a diagnostic diagram: high- radio galaxies (HzRGs) commonly have no detection of C iv (56 % in De Breuck et al. 2000) and SFGs usually show C iv in absorption instead of emission (e.g. Shapley et al. 2003; Calabrò et al. 2022). In Fig. 3 we show the line ratios of the candidate AGNs and SDSS quasars together with line ratios of SFGs and different types of AGNs from the literature. Both quasars and both AGN candidates as well as 30 % of the literature AGNs fall within the composite region of the diagram, indicating contribution from both star formation and AGN activity. Surprisingly, the line ratios of QSO1 indicate the highest contribution from star formation of all four objects although it is definitively classified as quasar, further highlighting the difficulty in AGN selection based on UV line ratios. It is also notable that the stacks of observed SFGs selected based on C iii] emission are classified as composite objects as well and show similar C3He2 ratios as AGNc1.
However, the detection of N v and the extreme line width of AGNc1 larger than firmly point towards an AGN origin. If these broad lines were produced by galactic-scale outflows instead, the violent kinematics should lead to a high contribution from shocks and likely lead to a detectable C iv emission line.
[…]
4 Discussion
4.1 Clues from the peculiar extended Ly emission
The pink contours in Fig. 1 show the 2 and 5 levels of extended Ly emission, corresponding to a surface brightness of and respectively, and determined after subtracting empirical quasar point spread functions and continuum sources from the combined cube. This analysis is described in detail in Herwig et al. (2024) and Lobos et al. (2023). The nebular emission extends for about 50 kpc in the south-east direction with respect to QSO1, starting from the projected location of AGNc2. Additional Ly emission coincides with the position of AGNc1. Surprisingly, the surface brightness level of the extended Ly emission is much lower than typically found for multiple AGN systems (Hennawi et al., 2015; Arrigoni Battaia et al., 2018) and the AGN candidates are not embedded in the nebula. Instead, the flux-weighted centroid of the extended emission is offset from QSO1 by 34 kpc (Herwig et al., 2024). Possible explanations for this could be a faster evolution and therefore early cool gas depletion of outlier density peaks in the universe or a later merger stage of the close AGN triplet candidate presented in this work. Moreover, strong feedback effects might have already depleted the circumgalactic medium of this system of cool gas.
The peculiar morphology of the extended Ly emission could be further evidence for violent on-going galaxy interactions: if the ISM of AGNc2 is ram-pressure stripped during the infall into the hot halo of QSO1, the trailing dense gas behind AGNc2 could become visible as extended Ly emission. Indeed, estimating the stripped gas mass yields a lower limit of at least when assuming the line-of-sight velocity, or roughly when using a conservative estimate of the 3D velocity, while the estimated cool gas mass visible as extended Ly emission amounts to roughly (Appendix C). This first estimate confirms that ram-pressure stripping could be an effective mechanism in this configuration and could indeed explain the peculiar appearance of the extended Ly emission. Evidence for similar ram-pressure stripping events around high-z quasars have been recently invoked by other studies on extended Lya emission (e.g., Chen et al. 2021).
4.2 Evolutionary scenario
The close separation of QSO1, AGNc1 and AGNc2 and evidence for gas stripping and therefore ongoing galaxy interactions makes it very likely that the host galaxies will quickly merge into one massive galaxy. The result of such a merger is predicted to have a host halo mass of (Bhowmick et al., 2020). Three-body interactions have a high likelihood of ejecting one of the participants via gravitational recoil (Partmann et al., 2024). This is the most probable evolution of the close black hole system, although two or even three mergers are also possible and would lead to the emission of gravitational waves possibly contributing to the gravitational wave background (Sesana et al., 2008) and likely the formation of an ultra-massive () black hole. However, ejection or stalling of one or two black holes is predicted to not hinder fast growth: in simulations, the strong gravitational forces lead to quick infall times of gas into the most massive black hole and facilitate black hole mass growth up to irrespective of coalescence (Hoffman et al., 2023). The extreme density peak implied by the presence of four AGNs in close proximity and three AGNs in the same dark matter halo is likely accompanied by a galaxy overdensity. This hypothesis needs to be tested by multi-wavelength observations targeting the galaxy population on Mpc scales around the quadruple AGN candidate. If confirmed, the most probable evolutionary path of this extraordinary system is the formation of a galaxy cluster in which the brightest cluster galaxy is formed out of the close triple AGN host galaxies.
4.3 Incidence rate of AGN tuples
Following Hennawi et al. (2015), an estimate of the probability to find such two close companion AGNs to a known quasar can be gained from the two-point correlation function for quasars and amounts to , or if considering any halo-scale triple AGN (Appendix D). A different generalized estimate for the AGN tuple incidence rate can be gained by considering the full observed parent sample. Including the candidate presented in this work, 2 AGN tuples have been identified within the QSO MUSEUM survey encompassing 134 quasars at observed with MUSE to similar depths (Arrigoni Battaia et al. 2019; Herwig et al. 2024; González Lobos et al. in preparation). This corresponds to an incidence rate of such systems of 1.5 % with Poisson errors for small-numbers statistics from Gehrels (1986), much larger than the above calculations and than previously believed (Hennawi et al., 2015). This implies undiscovered tuple systems around SDSS quasars (Pâris et al., 2018) in the same redshift range (with a median redshift of ), and, extrapolated to the entire sky, AGN tuples. Assuming that black hole tuples populate the massive end of the halo mass function and that 65 % of them are visible as AGN tuples (Hoffman et al., 2023), this number would amount to all halos down to M⊙ of the halo mass function (Behroozi et al., 2013).
5 Summary
We present an unlensed candidate quadruple AGN system at consisting of two SDSS type 1 quasars at a separation of roughly 480 kpc (QSO1 and QSO2) and two AGN candidates (AGNc1 and AGNc2) accompanying QSO1 in close proximity with a projected separation of up to 20 kpc. The identification as probable AGN candidates is based on C43 versus C4He2 line ratios, high ionization lines like N v in their spectrum and the emission line widths. AGNc1 is a potential type 1 AGN, while the spectral characteristics of AGNc2 align with those of a type 2 quasar. Further observations in X-ray or of rest-frame optical lines can confirm these objects as AGNs and allow for more reliable redshift determination.
The close triple system (QSO1, AGNc1, AGNc2) is associated with a Ly nebula, although it is not embedded within it. A possible explanation for the presence of peculiar Ly-bright gas is ram-pressure stripping of AGNc2 during its infall into the halo of QSO1.
This extraordinary system likely pinpoints a site of galaxy cluster formation. This hypothesis needs to be tested further by studying the Mpc-scale galaxy overdensity surrounding the quadruple AGN candidate.
This new candidate quadruple AGN allow us to refine the incidence rate of such rare systems around SDSS quasars to 1.5 %. Pre-selection of SDSS quasars with close faint sources (e.g., Euclid) is essential to build a statistical sample (12 systems), needed to confirm our statistics and constrain current models (e.g., Hoffman et al. 2023). Without it, additional MUSE observations (1 hour/source) of SDSS quasars would be required.
Acknowledgements.
We thank Guinevere Kauffmann for providing useful comments to this work. E.P.F. is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35
- Agnello et al. (2018) Agnello, A., Lin, H., Kuropatkin, N., et al. 2018, MNRAS, 479, 4345
- Arrigoni Battaia et al. (2022) Arrigoni Battaia, F., Chen, C.-C., Liu, H.-Y. B., et al. 2022, ApJ, 930, 72
- Arrigoni Battaia et al. (2019) Arrigoni Battaia, F., Hennawi, J. F., Prochaska, J. X., et al. 2019, MNRAS, 482, 3162
- Arrigoni Battaia et al. (2018) Arrigoni Battaia, F., Prochaska, J. X., Hennawi, J. F., et al. 2018, MNRAS, 473, 3907
- Assef et al. (2018) Assef, R. J., Stern, D., Noirot, G., et al. 2018, ApJS, 234, 23
- Bacon et al. (2010) Bacon, R., Accardo, M., Adjali, L., et al. 2010, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7735, Ground-based and Airborne Instrumentation for Astronomy III, ed. I. S. McLean, S. K. Ramsay, & H. Takami, 773508
- Becker et al. (1994) Becker, R. H., White, R. L., & Helfand, D. J. 1994, in Astronomical Society of the Pacific Conference Series, Vol. 61, Astronomical Data Analysis Software and Systems III, ed. D. R. Crabtree, R. J. Hanisch, & J. Barnes, 165
- Behroozi et al. (2013) Behroozi, P. S., Wechsler, R. H., & Conroy, C. 2013, ApJ, 770, 57
- Bhowmick et al. (2020) Bhowmick, A. K., Di Matteo, T., & Myers, A. D. 2020, MNRAS, 492, 5620
- Calabrò et al. (2022) Calabrò, A., Pentericci, L., Talia, M., et al. 2022, A&A, 667, A117
- Chen et al. (2021) Chen, C.-C., Arrigoni Battaia, F., Emonts, B. H. C., Lehnert, M. D., & Prochaska, J. X. 2021, ApJ, 923, 200
- De Breuck et al. (2000) De Breuck, C., Röttgering, H., Miley, G., van Breugel, W., & Best, P. 2000, A&A, 362, 519
- Djorgovski et al. (2007) Djorgovski, S. G., Courbin, F., Meylan, G., et al. 2007, ApJ, 662, L1
- Dullo et al. (2017) Dullo, B. T., Graham, A. W., & Knapen, J. H. 2017, MNRAS, 471, 2321
- Farina et al. (2019) Farina, E. P., Arrigoni-Battaia, F., Costa, T., et al. 2019, ApJ, 887, 196
- Farina et al. (2013) Farina, E. P., Montuori, C., Decarli, R., & Fumagalli, M. 2013, MNRAS, 431, 1019
- Feltre et al. (2016) Feltre, A., Charlot, S., & Gutkin, J. 2016, MNRAS, 456, 3354
- Fossati et al. (2021) Fossati, M., Fumagalli, M., Lofthouse, E. K., et al. 2021, MNRAS, 503, 3044
- Fukugita et al. (2004) Fukugita, M., Nakamura, O., Schneider, D. P., Doi, M., & Kashikawa, N. 2004, ApJ, 603, L65
- García-Vergara et al. (2022) García-Vergara, C., Rybak, M., Hodge, J., et al. 2022, ApJ, 927, 65
- Gehrels (1986) Gehrels, N. 1986, ApJ, 303, 336
- González Lobos et al. (2023) González Lobos, V., Arrigoni Battaia, F., Chang, S.-J., et al. 2023, A&A, 679, A41
- Green et al. (2011) Green, P. J., Myers, A. D., Barkhouse, W. A., et al. 2011, ApJ, 743, 81
- Gunn & Gott (1972) Gunn, J. E. & Gott, J. Richard, I. 1972, ApJ, 176, 1
- Guo et al. (2018) Guo, H., Shen, Y., & Wang, S. 2018, PyQSOFit: Python code to fit the spectrum of quasars, Astrophysics Source Code Library, record ascl:1809.008
- Habouzit et al. (2019) Habouzit, M., Volonteri, M., Somerville, R. S., et al. 2019, MNRAS, 489, 1206
- Hennawi et al. (2015) Hennawi, J. F., Prochaska, J. X., Cantalupo, S., & Arrigoni-Battaia, F. 2015, Science, 348, 779
- Hennawi et al. (2006) Hennawi, J. F., Strauss, M. A., Oguri, M., et al. 2006, AJ, 131, 1
- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., González Lobos, J., et al. 2024, arXiv e-prints, arXiv:2408.16826
- Hirschmann et al. (2019) Hirschmann, M., Charlot, S., Feltre, A., et al. 2019, MNRAS, 487, 333
- Hoffman et al. (2023) Hoffman, C., Chen, N., Di Matteo, T., et al. 2023, MNRAS, 524, 1987
- Husband et al. (2013) Husband, K., Bremer, M. N., Stanway, E. R., et al. 2013, MNRAS, 432, 2869
- Kalfountzou et al. (2017) Kalfountzou, E., Santos Lleo, M., & Trichas, M. 2017, ApJ, 851, L15
- Kayo & Oguri (2012) Kayo, I. & Oguri, M. 2012, MNRAS, 424, 1363
- Kellermann et al. (1989) Kellermann, K. I., Sramek, R., Schmidt, M., Shaffer, D. B., & Green, R. 1989, AJ, 98, 1195
- Kocevski et al. (2023) Kocevski, D. D., Barro, G., McGrath, E. J., et al. 2023, ApJ, 946, L14
- Lacy et al. (2020) Lacy, M., Baum, S. A., Chandler, C. J., et al. 2020, PASP, 132, 035001
- Lau et al. (2016) Lau, M. W., Prochaska, J. X., & Hennawi, J. F. 2016, ApJS, 226, 25
- Liu et al. (2011) Liu, X., Shen, Y., & Strauss, M. A. 2011, ApJ, 736, L7
- Llerena et al. (2022) Llerena, M., Amorín, R., Cullen, F., et al. 2022, A&A, 659, A16
- Lobos et al. (2023) Lobos, V. G., Arrigoni Battaia, F., Chang, S.-J., et al. 2023, A&A, 679, A41
- Matsuoka et al. (2009) Matsuoka, K., Nagao, T., Maiolino, R., Marconi, A., & Taniguchi, Y. 2009, A&A, 503, 721
- Mehrgan et al. (2019) Mehrgan, K., Thomas, J., Saglia, R., et al. 2019, ApJ, 887, 195
- Nagao et al. (2006) Nagao, T., Maiolino, R., & Marconi, A. 2006, A&A, 447, 863
- Nigoche-Netro et al. (2010) Nigoche-Netro, A., Aguerri, J. A. L., Lagos, P., et al. 2010, A&A, 516, A96
- Onoue et al. (2018) Onoue, M., Kashikawa, N., Uchiyama, H., et al. 2018, PASJ, 70, S31
- Overzier (2016) Overzier, R. A. 2016, A&A Rev., 24, 14
- Pâris et al. (2018) Pâris, I., Petitjean, P., Aubourg, É., et al. 2018, A&A, 613, A51
- Pâris et al. (2017) Pâris, I., Petitjean, P., Ross, N. P., et al. 2017, A&A, 597, A79
- Partmann et al. (2024) Partmann, C., Naab, T., Rantala, A., et al. 2024, MNRAS, 532, 4681
- Pfeifle et al. (2019) Pfeifle, R. W., Satyapal, S., Manzano-King, C., et al. 2019, ApJ, 883, 167
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Runnoe et al. (2012) Runnoe, J. C., Brotherton, M. S., & Shang, Z. 2012, MNRAS, 422, 478
- Sandrinelli et al. (2018) Sandrinelli, A., Falomo, R., Treves, A., Scarpa, R., & Uslenghi, M. 2018, MNRAS, 474, 4925
- Sesana et al. (2008) Sesana, A., Vecchio, A., & Colacino, C. N. 2008, MNRAS, 390, 192
- Shapley et al. (2003) Shapley, A. E., Steidel, C. C., Pettini, M., & Adelberger, K. L. 2003, ApJ, 588, 65
- Shen et al. (2020) Shen, X., Hopkins, P. F., Faucher-Giguère, C.-A., et al. 2020, MNRAS, 495, 3252
- Shen et al. (2016) Shen, Y., Brandt, W. N., Richards, G. T., et al. 2016, ApJ, 831, 7
- Shen et al. (2019) Shen, Y., Hall, P. B., Horne, K., et al. 2019, ApJS, 241, 34
- Soto et al. (2016) Soto, K. T., Lilly, S. J., Bacon, R., Richard, J., & Conseil, S. 2016, MNRAS, 458, 3210
- Trainor & Steidel (2012) Trainor, R. F. & Steidel, C. C. 2012, ApJ, 752, 39
- Uchiyama et al. (2018) Uchiyama, H., Toshikawa, J., Kashikawa, N., et al. 2018, PASJ, 70, S32
- Vestergaard & Peterson (2006) Vestergaard, M. & Peterson, B. M. 2006, ApJ, 641, 689
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Ward et al. (2024) Ward, E., de la Vega, A., Mobasher, B., et al. 2024, ApJ, 962, 176
- Weilbacher et al. (2020) Weilbacher, P. M., Palsa, R., Streicher, O., et al. 2020, A&A, 641, A28
- Weilbacher et al. (2016) Weilbacher, P. M., Streicher, O., & Palsa, R. 2016, MUSE-DRP: MUSE Data Reduction Pipeline, Astrophysics Source Code Library, record ascl:1610.004
- Weilbacher et al. (2014) Weilbacher, P. M., Streicher, O., Urrutia, T., et al. 2014, in Astronomical Society of the Pacific Conference Series, Vol. 485, Astronomical Data Analysis Software and Systems XXIII, ed. N. Manset & P. Forshay, 451
- White et al. (2012) White, M., Myers, A. D., Ross, N. P., et al. 2012, MNRAS, 424, 933
- Wijers & Schaye (2022) Wijers, N. A. & Schaye, J. 2022, MNRAS, 514, 5214
Appendix A Rejecting the possibility of gravitational lensing
A possible explanation for multiple quasars in close separation is gravitational lensing of a single source (e.g., Agnello et al. 2018). In this case, the spectra would be very similar with slight variations due to different light paths. While all AGN in this system appear to have significantly different spectra (Fig. 2), this view might be skewed by the low signal to noise ratio of the fainter objects. While AGNc2 is a type 2 AGN candidate as it only shows narrow emission lines and should therefore be viewed at an entirely different angle than QSO1 and AGNc1, we explore the possibility that the latter two objects are actually two lensed images of the same quasar.
In this case, we expect to find a massive object capable of acting as the gravitational lens in between the two images. To calculate the necessary magnitude of a lensing elliptical galaxy, we follow Farina et al. (2013), making use of the Faber–Jackson relation by Nigoche-Netro et al. (2010) and assuming an isothermal sphere to link mass and velocity dispersion. Assuming an Einstein ring radius of , corresponding to half of the angular distance between QSO1 and AGNc1, the minimal possible apparent -band magnitude of an interloper galaxy is at , brighter than both the AGN candidates and QSO1.
We then inspect the residuals of the point spread function modeling (Fig. 1, right). There, an artifact of the CCD gap becomes apparent as vertical line, but no foreground galaxy is revealed. The -band magnitude in a 1.6 aperture between QSO1 and AGNc1 (indicated as pink circle in Fig. 1, right) after masking negative values potentially arising due to the model subtraction is 26.6 mag, far below the faintest possible lensing galaxy. Therefore we conclude that no lensing geometry fits the observed pattern without revealing the lensing galaxy in the data cube.
Appendix B Spectral fitting
We attempt to fit a set of rest-frame UV lines in the spectra of all four sources, namely Ly, N v , C iv , He ii and C iii]. Due to the stark difference in magnitude and therefore data quality, we adopt two different methods. The two SDSS detected quasars QSO1 and QSO2 have sufficient S/N to be fitted using the package pyQSOFIT (Guo et al. 2018; Shen et al. 2019). We use the spectrum extracted from the MUSE cube as explained in Sec. 3.1 for QSO1 as it has higher S/N than the SDSS spectrum. For QSO2, only the SDSS spectrum is available. We smooth both by the same kernel size of pixels. Errors are estimated through 250 MCMC samplings, and the line width is determined from the fit to the C iii]line. Although we attempt to include a component for the N v line when fitting the Ly line complex of QSO2, we do not find a combination of priors that leads to convergence and we thus do not report a line flux for it. From pyQSOFIT we also obtain the continuum luminosity and the FWHM of C iv and use them to calculate black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived in Vestergaard & Peterson (2006). The obtained values are and .
Due to the low S/N in the continuum, the line spectra of candidates AGNc1 (smoothed by a 1D Gaussian kernel with pixels; Fig. 4) and AGNc2 (smoothed with pixels; Fig. 5) instead are fitted with Gaussian or double Gaussian line models and weighted by the noise spectra extracted from the rescaled variance cube. Fitting is performed using the function curve_fit in scipy (Virtanen et al. 2020).
Due to the highly non-Gaussian Ly line shape of AGNc1, the spectrum does not appear well-fit by the model. To test if this impacts the line flux, we smooth the spectrum further with a Gaussian kernel size of 10 pixels and repeat fitting. We find that the resulting line flux in both Ly and N v is decreases by less than 4 %. For the AGN candidates, the line widths are determined by calculating the second moment in an unsmoothed spectrum after subtracting the mean continuum level. The calculation window is chosen according to the width of the Gaussian line model fitted to the spectrum and centered on the Ly line peak (AGNc1) and the C iv line peak (AGNc2). Although these lines are resonant and therefore not ideal for line width determination, they have the highest S/N and are the only lines allowing for robust calculation of the second moment. Errors on are obtained by resampling the spectra with the associated noise spectra 10,000 times and, after calculating the second moment, the 25th and 75th percentiles of the obtained distribution are taken as lower and upper error.
A major source of uncertainty in the line flux of AGNc1 and AGNc2 is the poorly constrained underlying continuum. Especially the broad lines in the spectrum of AGNc1 are difficult to distinguish from the continuum and are affected by noise fluctuations, necessitating the large smoothing kernel size to be able to fit Gaussian profiles to the lines. Still, we can only place upper limits on C iv in the spectrum of AGNc1, and we also do not detect N v in the spectrum of AGNc2. Considering the large scatter in continuum level, we determine upper limits for the line fluxes by subtracting continuum implied by the minima closest to the expected line peak in the smoothed spectra and calculate moment 0 within 1 left and right of the expected line peak.
Appendix C Estimating the efficiency of ram pressure stripping
The peculiar extended Ly emission associated with the close triple system might be a further indication of on-going galaxy interaction as the emission could be a result of ram-pressure stripped ISM of AGNc2 appearing Ly bright. This could further explain the large velocity differences in the nebula obtained from the first moment with respect to AGNc2 (Fig. 6), reaching 1500 km s-1 at the East edge of the isophote furthest away from the AGN candidate and generally approaching the systemic redshift of AGNc2 closer to it, although the nebula shows significant turbulence. To determine the feasibility of this hypothesis, we estimate the efficiency of stripping in the configuration of the system.
Ram-pressure stripping is possible if (Gunn & Gott 1972). In this, is the density of the hot gas in the halo of QSO1 and obtained from the typical volume-weighted number density in a hot halo of mass log(/M at 20 % of the virial radius corresponding to the projected distance of AGNc2, (Wijers & Schaye 2022). This mass bin is chosen as it includes the halo mass estimate for single quasars at this redshift () as well as threefold this mass () in case QSO1, AGNc1 and AGNc2 all share one dark matter halo and all three resided in quasar-typical massive halos before. This larger halo mass is also similar to that estimated for one of the known triplets (Arrigoni Battaia et al. 2022). is the velocity with which AGNc2 is crossing the halo of QSO1 and at least as high as the line-of-sight velocity reported in Table 1. A conservative estimate of the 3D velocity is km s-1.
The gravitational pull of AGNc2 withstanding the ram pressure is described by the stellar and gas surface densities and . AGN hosts at typically have a massive (), compact stellar component described by a Sersic profile with an average Sersic index of 2.6 and an effective radius of 0.9 kpc (Kocevski et al. 2023). The gas disk is well-described by an exponential profile, and we assume an effective radius of 2 kpc (Ward et al. 2024) and a cold gas mass fraction of 20 %. The stripped gas mass can then be estimated by integrating the gas mass profile beyond the radius at which ram pressure exceeds the gravitational pull of the galaxy, and it amounts to at least when assuming the line-of-sight velocity, or roughly when using the 3D velocity.
To contrast this with the Ly-bright gas, we can estimate the mass from the extended emission using the equation with the Ly nebula area kpc2 (Herwig et al. 2024), the proton mass and hydrogen fraction and the typical CGM column density of cool gas, log( (Lau et al. 2016). The latter quantity is associated with the highest uncertainty as it can vary by one order of magnitude and it might not be applicable to extraordinary systems. With these values, we obtain an estimated cool gas mass in the Ly nebula of .
Appendix D Probability of finding a similarly close AGN triple
The probability to find two additional AGNs close to a known quasar can be estimated using
with the comoving number density of AGNs , the AGN-filled volume , the maximum distance between the AGNs and the two-point correlation function of quasars . This follows from a similar argument as presented in Hennawi et al. (2015), but we do not include the forth object, QSO2, in the estimation as the large projected distance implies that QSO2 does not share a host halo with the other three objects (yet) and thus, other assumptions have to be made in the calculation. Specifically, in the case of a triplet, equation 2 in Hennawi et al. (2015) will only contain three permutations for the two-point correlation function and one permutation for the three-point correlation function. Although the three-point correlation function relates to as (Hennawi et al. 2015), in our order-of-magnitude estimation, we can approximate , hence the factor of four in the above probability equation.
We infer from the quasar luminosity function at (Shen et al. 2020) with a lower magnitude limit of mag and the quasar bolometric luminosity correction from Runnoe et al. (2012) resulting in cMpc-3. Due to the imprecision in redshift determinations from broad QSO emission lines and the possibility of high peculiar velocities causing additional Doppler shifts, we extrapolate the maximum distance in the triplet, , as the highest projected distance plus a 50 % margin to account for projection effects. This yields a quasar-filled comoving volume of cMpc3. On small scales, the two-point correlation function is well described with with and cMpc (Kayo & Oguri 2012). Plugging in these values, we obtain .
The candidate triple AGN presented in this work is closer than previously found AGN triples. To get an estimate of the probability to find any AGN triple on halo scales, we repeat the calculation using kpc, corresponding roughly to the virial radius of a halo with mass , and obtain a probability of . In contrast, the probability to find three AGNs with kpc at random without considering clustering can be described with (Hennawi et al. 2015) and would be exceedingly rare with .

### v2

@@
 The extended Ly emission associated with the close triple system is more modest in extent and brightness compared to similar multiple AGN systems and could be caused by ram-pressure stripping of the type-2 quasar host during infall into the central dark matter halo.
-The projected evolution of the system into a galaxy cluster with the AGN host galaxies forming the brightest cluster galaxy needs to be further tested by galaxy overdensity studies on large scales around the quadruple AGN candidate.
+The predicted evolution of the system into a galaxy cluster with the AGN host galaxies forming the brightest cluster galaxy needs to be further tested by galaxy overdensity studies on large scales around the quadruple AGN candidate.
 If confirmed as a quadruple AGN with X-ray observations or rest-frame optical line ratios, this system would represent the second AGN quartet, the highest-redshift multiplet and the closest high-redshift triplet known.
 Key Words.:
-Galaxies: clusters: general – Galaxies: evolution – quasars: general1 Introduction
+Galaxies: interactions – Galaxies: high-redshift – Galaxies: evolution – quasars: general – quasars: emission lines1 Introduction
 In recent years, efforts to identify the progenitors of galaxy clusters, so called protoclusters (e.g., Overzier 2016), have accelerated.
@@
 While they often do reside in overdensities (Fossati et al., 2021; García-Vergara et al., 2022), their typical halo masses of at are much lower than cluster masses (White et al., 2012; Trainor & Steidel, 2012; Husband et al., 2013), requiring substantial growth to form a galaxy cluster until .
-While single bright quasars may pinpoint protocluster cores, physically associated multiple AGN systems could be even more reliable tracers of local overdensities (Onoue et al., 2018; Sandrinelli et al., 2018), albeit also not a firm indicator (Fukugita et al., 2004; Green et al., 2011), as a number of highly active, massive galaxies require ample supply of cold gas feeding supermassive black holes (SMBHs) and star formation.
+While single bright quasars may pinpoint protocluster cores, physically associated multiple active galactic nuclei (AGN) systems could be even more reliable tracers of local overdensities (Onoue et al., 2018; Sandrinelli et al., 2018), albeit also not a firm indicator (Fukugita et al., 2004; Green et al., 2011), as a number of highly active, massive galaxies require ample supply of cold gas feeding supermassive black holes (SMBHs) and star formation.
 In the framework of hierarchical structure formation, such objects are most likely to reside in peaks of the density field populating the rare massive end of the halo mass function (Bhowmick et al., 2020).
@@
 2019).
-So far, only a few multiple AGN systems on kiloparsec scales are known.
+So far, only a few high-z multiple AGN systems on kiloparsec scales are known.
 These objects are likely in an early stage of galaxy merger and might share a dark matter halo, but not their host galaxies.
@@
 2019).
-This work concentrates on the former AGN tuples on kiloparsec scales, of which only five are known.
+This work concentrates on the former AGN multiplets on kiloparsec scales, of which only five are known.
 Specifically, Hennawi et al.
@@
 Here, we report on the discovery of a quadruple AGN candidate at consisting of a close triplet, contained within 20 kpc in projection, and a forth quasar at a separation of 478 kpc.
-Throughout this work, we assume a flat CDM cosmology with , and (Planck Collaboration et al., 2020). Reported quasar magnitudes are in the AB system.
+Throughout this work, we assume a flat CDM cosmology with , and (Planck Collaboration et al., 2020).
+Reported quasar magnitudes are in the AB system.
 2 Observation and data analysis
@@
 For QSO2, only the SDSS spectrum is available.
+In SDSS DR17 (Abdurro’uf et al., 2022), the -band magnitude of the quasars integrated within the SDSS fiber with a diameter of 3 are mag and mag.
 Here we briefly describe the observations and data reduction.
@@
 Three exposures were taken with on-source time each, rotated by 90 degrees with respect to each other and covering a wavelength range of with a sampling of .
+At the wavelength of Ly, the spectral resolution of corresponds to km s-1 ( km s-1 at C iii]).
 The data have been reduced following Farina et al.
@@
 3.1 Discovery of a candidate quadruple AGN
-By collapsing the final MUSE cube along the wavelength axis, we obtain a continuum image of the field surrounding QSO1.
+By collapsing the final MUSE cube along the wavelength axis, we obtained a continuum image of the field surrounding QSO1.
 In this image (r-band in Fig.
 1, left), two additional continuum sources are visible in close proximity to QSO1.
-While similar geometries can be seen as a result of gravitational lensing of a single source, we can confidentily reject this possibility in this case (Appendix A).
+While similar geometries can be seen as a result of gravitational lensing of a single source, we can confidently reject this possibility in this case (Appendix A).
 We instead identify both of these sources (hereinafter called AGNc1 and AGNc2) as AGN candidates as detailed in the following.
-To constrain their nature, we extract spectra from the MUSE cube within an aperture of 0.8 to minimize contamination from each other’s PSFs.
-To determine their positions, we first model the PSF of a star within the MUSE FoV with a 2D Moffat function and fit the model to the three continuum sources observed with MUSE (QSO1, AGNc1, AGNc2) by fixing the FWHM and leaving amplitude and position free to vary (Fig.1, middle panel).
-The peak of the Moffat function is used as center of the aperture for spectral extraction and is indicated with pink circles (crosses) in the middle (right) panel of Fig.
+To constrain their nature, we extracted spectra from the MUSE cube within an aperture of 0.8 to minimize contamination from each other’s PSFs.
+To determine their positions, we first modeled the PSF of a star within the MUSE FoV with a 2D Moffat function and fit the model to the three continuum sources observed with MUSE (QSO1, AGNc1, AGNc2) by fixing the FWHM and leaving amplitude and position free to vary (Fig.
+1, middle panel).
+The peak of the Moffat function was used as center of the aperture for spectral extraction and is indicated with pink circles (crosses) in the middle (right) panel of Fig.
 1.
-To reduce noise while retaining the signal, the spectra are smoothed by a 1D Gaussian kernel with pixels, corresponding to .
-Before smoothing the spectrum of AGNc2, a sky emission line at 6365.3 is masked as it would otherwise be convolved with the C iv line.
+To reduce noise while retaining the signal, the spectra were smoothed by a 1D Gaussian kernel with pixels, corresponding to .
+Before smoothing the spectrum of AGNc2, we masked a sky emission line at 6365.3 as it would otherwise be convolved with the C iv line (Appendix E).
 The spectra are shown in Fig.
@@
 The two QSOs and two AGN candidates are all detected in Ly, confirming that they reside at roughly the same redshift.
-QSO1 and AGNc1 both show a broad C iii] emission line, while in the spectrum of AGNc2, there is a tentative detection of a narrow C iii] line as well as a prominent narrow, blue-shifted C iv line.
-A He ii emission line is tentatively detected in all four objects.
-Within the aperture of used for spectral extraction, the -band magnitude for the three AGNs observed with MUSE is mag, mag and mag.
-In SDSS DR17 (Abdurro’uf et al., 2022), the -band magnitude of QSO2 is reported as .
-However, the SDSS fiber diameter of 3 is much larger than the aperture used in this work and consequently, the magnitude of QSO1 is reported to be much brighter in SDSS with .
-Between QSO1 and AGNc1, we measure an angular separation of 2.1, corresponding to the distance kpc at this redshift.
+AGNc1 tentatively shows a broad C iii] line and has excess flux in N v , Si iv and He ii , while in the spectrum of AGNc2, there is a tentative detection of a narrow C iii] and He ii line as well as a prominent narrow, blue-shifted C iv line.
+Within the aperture of used for spectral extraction, the -band magnitude for the two AGN candidates is mag and mag111In this smaller aperture, the magnitude of QSO1 decreases to mag.
+Between QSO1 and AGNc1, we measured an angular separation of 2.1, corresponding to the distance kpc at this redshift.
 QSO1 and AGNc2 are separated by 2.6 or 20 kpc, and QSO2 is 61.6 or 478 kpc away from QSO1.
-We obtain the systemic redshift by fitting the C iii] line with a double Gaussian model and correcting by the characteristic velocity shift of between the peak of the line complex and the systemic redshift for quasars (Shen et al., 2016).
-The value obtained for AGNc1 is likely overestimated as a redshift of implies a highly unusual Ly line shape that is completely dominated by the blue wing (Fig.
+We obtained the systemic redshift by fitting the C iii] line with a double Gaussian model and using the peak of the line complex as reference wavelength.
+The value obtained for AGNc1 is likely overestimated as a redshift of implies a highly unusual Ly line shape dominated by the blue wing (Fig.
 4).
@@
 We fit the spectra as described in Appendix B and summarize the resulting line fluxes, widths and upper flux limits in Table 2.
-Black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived for C iv in Vestergaard & Peterson (2006) are and .
-The line width of AGNc2 is narrow with and consistent with a Type II AGN, while the spectrum of AGNc1 is dominated by broad lines of indicative of a Type I AGN.
+The line width of AGNc2 is narrow with km s-1 and consistent with a Type II AGN, while the spectrum of AGNc1 is dominated by broad lines of km s-1 indicative of a Type I AGN.
+3.2 Properties of QSO1 and QSO2
+Using the fitting results (Appendix B), we can derive additional properties for the brighter SDSS quasars QSO1 and QSO2.
+We calculated the black hole masses using the relation and absolute accuracy estimate derived for C iv in Vestergaard & Peterson (2006) and obtained and .
 At the position of QSO1, radio emission is detected with a flux density of in VLASS (Lacy et al., 2020) and in VLA FIRST (Becker et al., 1994), leading to a radio slope of ().
-Additionally using the UV continuum slope from the fit to the spectrum of QSO1 (Appendix B) of (), we can calculate the ratio (Kellermann et al., 1989) by extrapolating the rest-frame flux densities and obtain , classifying QSO1 as radio-loud.
-Although no additional radio emission is detected, we cannot exclude that AGNc1 and AGNc2 are radio-loud due to the shallow surveys depth and their dim UV continuum.
+Additionally using the UV continuum slope from the fit to the spectrum of QSO1 (Appendix B) of (), we calculated the ratio (Kellermann et al., 1989) by extrapolating the rest-frame flux densities and obtained , classifying QSO1 as radio-loud.
+Although no additional radio emission is detected, we cannot exclude that AGNc1 and AGNc2 are radio-loud due to the shallow surveys depth222The VLASS continuum image rms is Jy/beam (Lacy et al., 2020), FIRST data has an rms of Jy/beam close to the system location(Becker et al., 1994).
+and their dim UV continuum.
 […]
-QSO1 | 10:12:54.74 | +03:35:48.93 | 3.166 0.001 | 0 | 0/ 0 | |
-AGNc1 | 10:12:54.88 | +03:35:49.22 | 3.183 0.008 | -1590 766 | 2.1/16 | |
-AGNc2 | 10:12:54.77 | +03:35:46.36 | 3.158 0.001 | 761 166 | 2.6/20 | |
 […]
-3.1.1 Emission line diagnostics
+| AGNc1 | 10:12:54.88 | +03:35:49.22 | 3.181 0.008 | -1590 766 | 2.1/16 | |
+| AGNc2 | 10:12:54.77 | +03:35:46.36 | 3.156 0.001 | 761 166 | 2.6/20 | |
+[…]
+3.3 Emission line diagnostics
 A commonly adopted diagnostic diagram to distinguish AGNs from star forming galaxies (SFGs) in rest-frame UV uses the line ratios C iv /C iii] (C43), mainly influenced by the ionization parameter, and C iii]/He ii (C3He2), which is sensitive to metallicity and allows distinction between shocks and AGN photoionization (Feltre et al., 2016).
@@
 Classification is further complicated because certain types of galaxies would not appear in such a diagnostic diagram: high- radio galaxies (HzRGs) commonly have no detection of C iv (56 % in De Breuck et al.
-2000) and SFGs usually show C iv in absorption instead of emission (e.g.
+2000) and SFGs usually show C iv and Si iv in absorption instead of emission (e.g.
 Shapley et al.
@@
 Surprisingly, the line ratios of QSO1 indicate the highest contribution from star formation of all four objects although it is definitively classified as quasar, further highlighting the difficulty in AGN selection based on UV line ratios.
-It is also notable that the stacks of observed SFGs selected based on C iii] emission are classified as composite objects as well and show similar C3He2 ratios as AGNc1.
-However, the detection of N v and the extreme line width of AGNc1 larger than firmly point towards an AGN origin.
+It is also notable that the stacks of observed SFGs selected based on C iii] emission are classified as composite objects as well.
+However, the detection of N v and Si iv , two highly ionized lines, and the extreme line width of AGNc1 larger than firmly point toward an AGN origin.
 If these broad lines were produced by galactic-scale outflows instead, the violent kinematics should lead to a high contribution from shocks and likely lead to a detectable C iv emission line.
-[…]
-[…]
-[…]
-[…]
-[…]
+The upper limit on the EW of C iv lies above the AGN demarcation line of at the corresponding C iv /He ii ratio limit for AGNc1 of ¡ 1 (Nakajima et al., 2018) and AGN ionization can therefore not be excluded based on the non-detection of the line.
+The ratio of Si iv to He ii exceeds those typical for HzRGs (Humphrey et al., 2008) as well as quasars (Nagao et al., 2006).
+On the other hand, the C iv EW of AGNc2 lies firmly above the demarcation line of 12 Å in Nakajima et al.
+(2018) at the corresponding C iv /He ii ratio for AGNc2, indicating that this source is an AGN.
 4 Discussion
@@
 This analysis is described in detail in Herwig et al.
-(2024) and Lobos et al.
+(2024) and González Lobos et al.
 (2023).
@@
 If confirmed, the most probable evolutionary path of this extraordinary system is the formation of a galaxy cluster in which the brightest cluster galaxy is formed out of the close triple AGN host galaxies.
-4.3 Incidence rate of AGN tuples
+4.3 Incidence rate of AGN multiplets
 Following Hennawi et al.
-(2015), an estimate of the probability to find such two close companion AGNs to a known quasar can be gained from the two-point correlation function for quasars and amounts to , or if considering any halo-scale triple AGN (Appendix D).
-A different generalized estimate for the AGN tuple incidence rate can be gained by considering the full observed parent sample.
-Including the candidate presented in this work, 2 AGN tuples have been identified within the QSO MUSEUM survey encompassing 134 quasars at observed with MUSE to similar depths (Arrigoni Battaia et al.
+(2015), an estimate of the probability to find such two close companion AGNs to a known quasar can be gained from the two-point correlation function for quasars and, for a magnitude limit of 25 mag, amounts to , or if considering any halo-scale triple AGN (Appendix D).
+A different generalized estimate for the AGN multiplet incidence rate can be gained by considering the full observed parent sample.
+Including the candidate presented in this work, 2 AGN multiplets have been identified within the QSO MUSEUM survey encompassing 134 quasars at observed with MUSE to similar depths (Arrigoni Battaia et al.
 2019; Herwig et al.
@@
 This corresponds to an incidence rate of such systems of 1.5 % with Poisson errors for small-numbers statistics from Gehrels (1986), much larger than the above calculations and than previously believed (Hennawi et al., 2015).
-This implies undiscovered tuple systems around SDSS quasars (Pâris et al., 2018) in the same redshift range (with a median redshift of ), and, extrapolated to the entire sky, AGN tuples.
-Assuming that black hole tuples populate the massive end of the halo mass function and that 65 % of them are visible as AGN tuples (Hoffman et al., 2023), this number would amount to all halos down to M⊙ of the halo mass function (Behroozi et al., 2013).
+This implies undiscovered multiplet systems around SDSS quasars (Pâris et al., 2018) in the same redshift range (with a median redshift of ), and, extrapolated to the entire sky, AGN multiplet.
+Assuming that black hole multiplets populate the massive end of the halo mass function and that 65 % of them are visible as AGN multiplets (Hoffman et al., 2023), this number would amount to all halos down to M⊙ of the halo mass function (Behroozi et al., 2013).
 5 Summary
@@
 Acknowledgements.
-We thank Guinevere Kauffmann for providing useful comments to this work.
+We thank the anonymous referee for their constructive comments that helped improve the manuscript and Guinevere Kauffmann for providing useful comments to an earlier version of this work.
 E.P.F.
@@
 - Hennawi et al. (2006) Hennawi, J. F., Strauss, M. A., Oguri, M., et al. 2006, AJ, 131, 1
-- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., González Lobos, J., et al. 2024, arXiv e-prints, arXiv:2408.16826
+- Herwig et al. (2024) Herwig, E., Arrigoni Battaia, F., González Lobos, J., et al. 2024, A&A, 691, A210
 - Hirschmann et al. (2019) Hirschmann, M., Charlot, S., Feltre, A., et al. 2019, MNRAS, 487, 333
 - Hoffman et al. (2023) Hoffman, C., Chen, N., Di Matteo, T., et al. 2023, MNRAS, 524, 1987
+- Humphrey et al. (2008) Humphrey, A., Villar-Martín, M., Vernet, J., et al. 2008, MNRAS, 383, 11
 - Husband et al. (2013) Husband, K., Bremer, M. N., Stanway, E. R., et al. 2013, MNRAS, 432, 2869
@@
 - Llerena et al. (2022) Llerena, M., Amorín, R., Cullen, F., et al. 2022, A&A, 659, A16
-- Lobos et al. (2023) Lobos, V. G., Arrigoni Battaia, F., Chang, S.-J., et al. 2023, A&A, 679, A41
 - Matsuoka et al. (2009) Matsuoka, K., Nagao, T., Maiolino, R., Marconi, A., & Taniguchi, Y. 2009, A&A, 503, 721
@@
 - Nagao et al. (2006) Nagao, T., Maiolino, R., & Marconi, A. 2006, A&A, 447, 863
+- Nakajima et al. (2018) Nakajima, K., Schaerer, D., Le Fèvre, O., et al. 2018, A&A, 612, A94
 - Nigoche-Netro et al. (2010) Nigoche-Netro, A., Aguerri, J. A. L., Lagos, P., et al. 2010, A&A, 516, A96
@@
 - Shen et al. (2020) Shen, X., Hopkins, P. F., Faucher-Giguère, C.-A., et al. 2020, MNRAS, 495, 3252
-- Shen et al. (2016) Shen, Y., Brandt, W. N., Richards, G. T., et al. 2016, ApJ, 831, 7
 - Shen et al. (2019) Shen, Y., Hall, P. B., Horne, K., et al. 2019, ApJS, 241, 34
@@
 In this case, we expect to find a massive object capable of acting as the gravitational lens in between the two images.
-To calculate the necessary magnitude of a lensing elliptical galaxy, we follow Farina et al.
+To calculate the necessary magnitude of a lensing elliptical galaxy, we followed Farina et al.
 (2013), making use of the Faber–Jackson relation by Nigoche-Netro et al.
@@
 Assuming an Einstein ring radius of , corresponding to half of the angular distance between QSO1 and AGNc1, the minimal possible apparent -band magnitude of an interloper galaxy is at , brighter than both the AGN candidates and QSO1.
-We then inspect the residuals of the point spread function modeling (Fig.
+We then inspected the residuals of the point spread function modeling (Fig.
 1, right).
@@
 The -band magnitude in a 1.6 aperture between QSO1 and AGNc1 (indicated as pink circle in Fig.
-1, right) after masking negative values potentially arising due to the model subtraction is 26.6 mag, far below the faintest possible lensing galaxy.
+1, right) after masking negative values potentially arising due to the model subtraction is 26.4 mag, far below the faintest possible lensing galaxy.
 Therefore we conclude that no lensing geometry fits the observed pattern without revealing the lensing galaxy in the data cube.
 Appendix B Spectral fitting
-We attempt to fit a set of rest-frame UV lines in the spectra of all four sources, namely Ly, N v , C iv , He ii and C iii].
-Due to the stark difference in magnitude and therefore data quality, we adopt two different methods.
+We attempted to fit a set of rest-frame UV lines to the spectra of all four sources presented in Fig.
+2, namely Ly, N v , Si iv , C iv , He ii and C iii].
+Due to the stark difference in magnitude and therefore data quality, we adopted two different methods.
+All line fluxes, equivalent widths (EWs) and line widths are summarized in Table 2.
+[…]
+[…]
+[…]
+[…]
+[…]
+B.1 SDSS QSOs
 The two SDSS detected quasars QSO1 and QSO2 have sufficient S/N to be fitted using the package pyQSOFIT (Guo et al.
@@
 2019).
-We use the spectrum extracted from the MUSE cube as explained in Sec.
+We used the spectrum extracted from the MUSE cube as explained in Sec.
 3.1 for QSO1 as it has higher S/N than the SDSS spectrum.
 For QSO2, only the SDSS spectrum is available.
-We smooth both by the same kernel size of pixels.
-Errors are estimated through 250 MCMC samplings, and the line width is determined from the fit to the C iii]line.
-Although we attempt to include a component for the N v line when fitting the Ly line complex of QSO2, we do not find a combination of priors that leads to convergence and we thus do not report a line flux for it.
-From pyQSOFIT we also obtain the continuum luminosity and the FWHM of C iv and use them to calculate black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived in Vestergaard & Peterson (2006).
-The obtained values are and .
-Due to the low S/N in the continuum, the line spectra of candidates AGNc1 (smoothed by a 1D Gaussian kernel with pixels; Fig.
-4) and AGNc2 (smoothed with pixels; Fig.
-5) instead are fitted with Gaussian or double Gaussian line models and weighted by the noise spectra extracted from the rescaled variance cube.
-Fitting is performed using the function curve_fit in scipy (Virtanen et al.
+Errors were estimated through 250 MCMC samplings, and the line width was determined from the fit to the C iii] line.
+The Gaussian model used to describe the C iii] line is affected by the Gaussian smoothing kernel applied before fitting, and we thus corrected the values down to reflect the intrinsic line width.
+Although we attempted to include a component for the N v and Si iv line when fitting QSO2, we did not find a combination of priors that leads to convergence and we thus do not report fluxes for these lines.
+From pyQSOFIT we also obtained the continuum luminosity and the FWHM of C iv and used them to calculate black hole mass estimates for QSO1 and QSO2 using the relation and absolute accuracy estimate derived in Vestergaard & Peterson (2006).
+With erg s-1, erg s-1, km s-1 and km s-1, we obtained the values and .
+B.2 AGN candidates
+Due to low S/N, the spectra of candidates AGNc1 and AGNc2 cannot be fitted with pyQSOFIT and we instead subtracted a constant local continuum value before fitting emission lines using the function curve_fit in scipy (Virtanen et al.
 2020).
-Due to the highly non-Gaussian Ly line shape of AGNc1, the spectrum does not appear well-fit by the model.
-To test if this impacts the line flux, we smooth the spectrum further with a Gaussian kernel size of 10 pixels and repeat fitting.
-We find that the resulting line flux in both Ly and N v is decreases by less than 4 %.
-For the AGN candidates, the line widths are determined by calculating the second moment in an unsmoothed spectrum after subtracting the mean continuum level.
-The calculation window is chosen according to the width of the Gaussian line model fitted to the spectrum and centered on the Ly line peak (AGNc1) and the C iv line peak (AGNc2).
-Although these lines are resonant and therefore not ideal for line width determination, they have the highest S/N and are the only lines allowing for robust calculation of the second moment.
-Errors on are obtained by resampling the spectra with the associated noise spectra 10,000 times and, after calculating the second moment, the 25th and 75th percentiles of the obtained distribution are taken as lower and upper error.
-A major source of uncertainty in the line flux of AGNc1 and AGNc2 is the poorly constrained underlying continuum.
-Especially the broad lines in the spectrum of AGNc1 are difficult to distinguish from the continuum and are affected by noise fluctuations, necessitating the large smoothing kernel size to be able to fit Gaussian profiles to the lines.
-Still, we can only place upper limits on C iv in the spectrum of AGNc1, and we also do not detect N v in the spectrum of AGNc2.
-Considering the large scatter in continuum level, we determine upper limits for the line fluxes by subtracting continuum implied by the minima closest to the expected line peak in the smoothed spectra and calculate moment 0 within 1 left and right of the expected line peak.
+In particular, we obtained a first estimate of the line widths by calculating the second moment of the Ly line (AGNc1) and C iv line (AGNc2) in an unsmoothed spectrum.
+We used the resulting values of km s-1 and km s-1 respectively (corresponding to and respectively) to constrain the following fit.
+Then, we calculated the continuum level and its variance by masking all shaded gray areas in Fig.
+2 as well as marked emission lines within (AGNc1) and (AGNc2) of the systemic wavelength.
+The chosen masking windows are large enough to ensure no line emission is contaminating the continuum calculation and small enough to retain local continuum around each fitted line.
+The latter is particularly important for the broad line width of AGNc1.
+Due to the large velocity shift of C iv in AGNc2, we adjusted the wavelength of the line downward by to ensure appropriate masking and fitting.
+For each fitted emission line, we then considered the next 40 unmasked wavelength channels (corresponding to 50 Å) left and right of the expected peak and calculated the 25th, 50th, and 75th percentile to get the median continuum value (Cont50) and typical range (Cont25, Cont75).
+Cont25 is consistent with the noise level for most lines (Fig.
+4, 5) and thus represents a conservative lower limit for the continuum level as the AGN candidates are detected in continuum emission (Fig.
+1).
+As the determination of the continuum introduces the largest uncertainty in the fit, in particular for the broad lines of AGNc1, we performed three emission line fittings using each of these continuum values for the subtraction and derived conservative line flux errors from the difference in results.
+During line fitting, we assumed a single Gaussian model within the bounds given in Table 3, and weighted the fit with noise spectra extracted from the rescaled variance cube.
+The Ly and N v line of AGNc1 are likely blended together and are thus fitted together using two Gaussian lines centered on the respective systemic wavelengths.
+From the Gaussian model of the C iii] line, we obtained the line width of km s-1 (AGNc1) and km s-1 (AGNc2).
+These values are larger than the calculation of the second moment in unsmoothed spectra, but are consistent with line widths in type 1 AGNs and type 2 AGNs respectively.
+We again corrected the line width of the Gaussian model by the smoothing kernel.
+| Object | Normalization | Peak | |
+| AGNc1 | (0, inf) | (, ) | |
+| AGNc2 | (0, inf) | (, ) |
+If the flux value of an emission line is consistent with zero after subtracting Cont75, we consider it as undetected and only place upper limits on the flux.
+This is the case for C iv (AGNc1) and N v and Si iv (AGNc2).
+We derived upper limits from the channel-wise RMS of the spectra scaled to and report the limit at significance.
+We derived EWs and uncertainties for the high ionization lines Si iv and C iv from the three different Gaussian line fittings and their respective continuum levels and derived upper limits on EWs by assuming the median continuum level Cont50 and the upper flux limit.
 Appendix C Estimating the efficiency of ram pressure stripping
@@
 6), reaching 1500 km s-1 at the East edge of the isophote furthest away from the AGN candidate and generally approaching the systemic redshift of AGNc2 closer to it, although the nebula shows significant turbulence.
-To determine the feasibility of this hypothesis, we estimate the efficiency of stripping in the configuration of the system.
+To determine the feasibility of this hypothesis, we estimated the efficiency of stripping in the configuration of the system.
 Ram-pressure stripping is possible if (Gunn & Gott 1972).
@@
 2023).
-The gas disk is well-described by an exponential profile, and we assume an effective radius of 2 kpc (Ward et al.
+The gas disk is well-described by an exponential profile, and we assumed an effective radius of 2 kpc (Ward et al.
 2024) and a cold gas mass fraction of 20 %.
 The stripped gas mass can then be estimated by integrating the gas mass profile beyond the radius at which ram pressure exceeds the gravitational pull of the galaxy, and it amounts to at least when assuming the line-of-sight velocity, or roughly when using the 3D velocity.
-To contrast this with the Ly-bright gas, we can estimate the mass from the extended emission using the equation with the Ly nebula area kpc2 (Herwig et al.
-2024), the proton mass and hydrogen fraction and the typical CGM column density of cool gas, log( (Lau et al.
+To contrast this with the Ly-bright gas, we estimated the mass from the extended emission using the equation with the Ly nebula area kpc2 (Herwig et al.
+2024), the proton mass and hydrogen fraction and the typical column density of cool gas in the circumgalactic medium, log( (Lau et al.
 2016).
 The latter quantity is associated with the highest uncertainty as it can vary by one order of magnitude and it might not be applicable to extraordinary systems.
-With these values, we obtain an estimated cool gas mass in the Ly nebula of .
-Appendix D Probability of finding a similarly close AGN triple
+With these values, we obtained an estimated cool gas mass in the Ly nebula of .
+Appendix D Probability of finding a similarly close AGN triplet
 The probability to find two additional AGNs close to a known quasar can be estimated using
@@
 2015), in our order-of-magnitude estimation, we can approximate , hence the factor of four in the above probability equation.
-We infer from the quasar luminosity function at (Shen et al.
+We inferred from the quasar luminosity function at (Shen et al.
 2020) with a lower magnitude limit of mag and the quasar bolometric luminosity correction from Runnoe et al.
 (2012) resulting in cMpc-3.
-Due to the imprecision in redshift determinations from broad QSO emission lines and the possibility of high peculiar velocities causing additional Doppler shifts, we extrapolate the maximum distance in the triplet, , as the highest projected distance plus a 50 % margin to account for projection effects.
+Due to the imprecision in redshift determinations from broad QSO emission lines and the possibility of high peculiar velocities causing additional Doppler shifts, we extrapolated the maximum distance in the triplet, , as the highest projected distance plus a 50 % margin to account for projection effects.
 This yields a quasar-filled comoving volume of cMpc3.
 On small scales, the two-point correlation function is well described with with and cMpc (Kayo & Oguri 2012).
-Plugging in these values, we obtain .
+Plugging in these values, we obtained .
 The candidate triple AGN presented in this work is closer than previously found AGN triples.
-To get an estimate of the probability to find any AGN triple on halo scales, we repeat the calculation using kpc, corresponding roughly to the virial radius of a halo with mass , and obtain a probability of .
+To get an estimate of the probability to find any AGN triple on halo scales, we repeated the calculation using kpc, corresponding roughly to the virial radius of a halo with mass , and obtained a probability of .
 In contrast, the probability to find three AGNs with kpc at random without considering clustering can be described with (Hennawi et al.
 2015) and would be exceedingly rare with .
+Appendix E Examining possible sources of spectral contamination
+Multiple sources of artifacts and spurious signal can contaminate observational data. Here, we examine a few of these effects and determine the level of contamination introduced.
+The first artifact to consider is residual emission from the CCD gap visible in the continuum image after subtraction of QSO1, AGNc1 and AGNc2 (Fig.
+1, right panel).
+In particular, the spectrum of AGNc2 overlaps with the CCD gap.
+Thus, we extracted a spectrum at the same RA as AGNc2, but close to a flux peak on the CCD gap, and rescaled it to match the mean pixel value at the position of AGNc2 in the residual image.
+This region is indicated by a white plus in Fig.
+1.
+The spectrum, smoothed with the same Gaussian kernel of size pixels, is displayed in Fig.
+7.
+For reference, we also show the spectrum of AGNc2 as well as the expected wavelengths of emission lines.
+While the CCD gap might introduce a noisy contribution to the continuum, it does not show strong emission features coinciding with the fitted emission lines of AGNc2.
+Next, we examined the level of background emission in the cube by visually selecting ten regions within the same aperture as the spectra with a radius of 0.8 without continuum emission, taking the mean of the extracted spectra and smoothing it with the Gaussian kernel of size pixels.
+The obtained spectrum is also displayed in Fig.
+7 and represents the typical channel-wise background level in the data.
+Due to its low level, it is not expected to contribute significantly to the continuum emission in any part of the spectra.
+A third source of contamination are strong atmospheric emission lines contributing significant flux peaks for the faint sources considered in this work.
+Despite careful subtraction of the sky spectrum in two different steps during the data reduction and before combining the individual exposures, residuals of lines can still remain in the combined data (Soto et al.
+2016).
+These residuals appear as very thin (one to three pixel wide) emission lines in the spectrum of continuum sources if they were masked together with source emission lines during sky removal.
+The spectrum of AGNc2 is specifically affected by this as there are multiple narrow lines visible in the unsmoothed spectrum of the combined cube as well as the individual exposures.
+In Fig.
+8, we show the affected spectral windows together with a sky emission model555Generated with the ”The Cerro Paranal Advanced Sky Model”, https://www.eso.org/observing/etc/doc/skycalc/The_Cerro_Paranal_Advanced_Sky_Model.pdf.
+In these unsmoothed spectra, the difference in line width to C iv becomes apparent.
+Each of the narrow emission features not assignable to a common AGN UV line coincides with the position of a sky emission line within one or two pixels.
+Therefore, these additional lines are most likely sky subtraction residuals.
+Finally, given the small number of exposures obtained, individual datacubes could contribute significantly to emission lines, particularly if the line width is narrow.
+To exclude the possibility that the narrow lines in the spectrum of AGNc2 are spurious, we excluded in turn one of the three exposures during combination of the final cube and examined the Ly and C iv lines in the different resulting spectra in comparison to the spectrum presented in Fig.
+2.
+As shown in Fig.
+9, the line emission persists irrespective of the combination of exposures.
+We caution that the gap between CCDs, which is masked before combination of the cubes, covers parts of the PSF of AGNc2 in the third exposure.
+However, when combining only two cubes, masked areas in one data cube are filled in with the values of the other data cube.
+This is the reason why the seemingly decreased line flux of exposure 3 (Fig.
+8) does not have a strong affect on the line flux in the combined cube (Fig.
+9).